OpenStudy (anonymous):

\left(\begin{matrix}sin ^{2}theta \\ 1-cos \left( theta \right)\end{matrix}\right). Trying to simplify this...having troublee :(

OpenStudy (anonymous):

I fail at writing notations

OpenStudy (anonymous):

^^the attachment is what i was trying to type

OpenStudy (callisto):

Don't know if it is really 'simplify' :| \[\frac{sin^2\theta}{1-cos\theta} = \frac{1-cos^2\theta}{1-cos\theta} = \frac{(1-cos\theta)(1+cos\theta)}{1-cos\theta} = 1+cos\theta\]

OpenStudy (anonymous):

Thats what I got! but they want me to select one of 4 and 1+cos(theta) is not one of them :(

OpenStudy (callisto):

Do you mind showing your options?

OpenStudy (callisto):

I mean all the options

OpenStudy (anonymous):

yes one sec

OpenStudy (anonymous):

i hope it works

OpenStudy (callisto):

Actually, you can simplify the options from the choices so that you can get the expression you've simplified.. for example, for (a) \[\frac{1+sec\theta}{sec\theta} = \frac{1}{sec\theta} + \frac{sec\theta}{sec\theta} = cos\theta +1\]

OpenStudy (anonymous):

ohhh cool! THanksssssssss :D I never would have figured that out!

OpenStudy (callisto):

Try the do something for the options, then you can discover the answer!~